OpenStudy (anonymous):

Find sin θ if cot θ = - 4 and cos θ < 0.

jhonyy9 (jhonyy9):

us cot x =cos x / sin x

OpenStudy (anonymous):

I have, and I understand that sin(x) must be positive. sin(x) must equal cos(x)/-4 right?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

I don't know what to do next. The multiple choice answers are: -17, sqrt(17)/17, -1/4, and sqrt(17)/4

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

\[\frac{ \cos \theta }{ \sin \theta } = -4\]\[\cos \theta = -4\sin \theta\]You can substitute that expression for cos in the following:\[\cos ^{2}\theta + \sin ^{2}\theta = 1\]

OpenStudy (anonymous):

\[\left( -4\sin \theta \right)^{2} + \sin ^{2}\theta = 1\]

OpenStudy (anonymous):

\[16\sin ^{2}\theta + \sin ^{2}\theta = 1\]From here, you should be able to simplify and get the answer.

OpenStudy (anonymous):

When simplifying, as a last step, rationalize your denominator.

OpenStudy (anonymous):

Also, bear in mind that the cotangent and cosine are negative. Therefore, the sine is positive.

OpenStudy (dumbcow):

another approach is to use a right triangle cot = adj/opp  the hypotenuse can be found using pythagorean thm

OpenStudy (anonymous):

Thanks guys!!! Really appreciate it.